Two balls with masses of of 2.5 kg and 6.2 kg travel toward each other at speeds of 9 m/s and 3.5 m/s, respectively. If the balls have a head-on inelastic collision and the 2.5-kilogram ball recoils with a speed of 7.00 m/s, how much kinetic energy is lost in the collision?
Answer #1

m1u1 + m2u2 = (m1v1 + m2v2)

2.5*9- 6.2*3.4 = (-2.5*7 + 6.2V)

=> V = 2.17 m/s

KE loss = KE1 - KE2 = (1/2*2.5*(9)^2 + 1/2*6.2*(3.5)^2 )- (1/2*2.5*(7)^2 + 1/2*6.2*(2.17)^2) = 63.37 J